A pair of balanced dice is tossed. If X equals the total number of spots showing...

  1. A pair of balanced dice is tossed. If X equals the total number of spots showing on the

dice, then, for k = 2, 3,...,12, find:

a. Pr(4≤X≤9)

b. Pr(4<X≤9)

c. Pr(4≤X<9)

d. Pr(4<X<9)
Answer #1

a)

cdf of sum x

x n p cdf
2 1 0.027778 0.027778
3 2 0.055556 0.083333
4 3 0.083333 0.166667
5 4 0.111111 0.277778
6 5 0.138889 0.416667
7 6 0.166667 0.583333
8 5 0.138889 0.722222
9 4 0.111111 0.833333
10 3 0.083333 0.916667
11 2 0.055556 0.972222
12 1 0.027778 1

P(4<= X <= 9)

= P(X<= 9) - P(X<= 3)

= 0.83333 - 0.08333

= 0.75

b)

P(4 < X<= 9)

= P(X<= 9) - P(X<= 4)

= 0.83333 - 0.166666

= 0.66666

c. Pr(4≤X<9)

= P(X<= 8) - P(X<= 3)

= 0.72222- 0.08333

= 0.63889

d. Pr(4<X<9)

= P(X<= 8) - P(X<= 4)

=0.72222 - 0.166666

= 0.555555